Consider a recurrence relation an = -3an-1 + n for n = 1,2,3,4,… with initial conditions a1 = 3. Calculate a3.

1
Expert's answer
2021-11-25T23:44:45-0500

Recurrence relation:

"a_{n} = -3a_{n-1} + n"

Initial condition:

"a_{1} = 3"


Use recurrence relation for "n = 2" :

"a_{2} = -3a_{1} + 2 = -3 * 3 + 2 =-9 + 2 = -7"


Use recurrence relation for "n = 3" :

"a_{3} = -3a_{2} + 3 = -3 * (-7) + 3 = 21 + 3 = 24"


Answer: "a_{3} = 24"